OpenStudy (anonymous):

Currently, Sam's coach is three times older than Sam. When Sam started soccer 5 years ago, his coach was five times his age. How old is Sam now?

OpenStudy (anonymous):

let sam=x let the coach= 3x x-5=3x(5) solve from there

OpenStudy (aakashsudhakar):

Sam's current age = S Coach's current age = C C = 3S [Sam's coach is three times older than Sam] 5(S-5) + 5 = C So you have a system of equations. C = 3S C = 5(S - 5) + 5 Solve for any variable of your choosing. I'm going to solve for S because I think it's easier. [C = 3S] [3S] = 5(S - 5) + 5 3S = 5S - 25 + 5 3S = 5S - 20 -2S = -20 S = 10 We've found that Sam's current age is 10 years old. Given that the coach is currently three times Sam's current age, the rest of this problem is very easy. 3S = C 3(10) = C C = 30 We've found that Sam's coach's current age if 30 years old. It's always important to cross-check these results to see if the results are accurate. Use the other equation in the system of equations to verify. C = 5(S - 5) + 5 30 = 5(10 - 5) + 5 30 = 5(5) + 5 30 = 25 + 5 30 = 30 The answers we've gathered are accurate. Sam's current age is 10 years old while his coach's current age is 30 years old.